I always have trouble with fractions, Im not sure why they put the 5/2 in parenthese,
3/7 x (5/2) =

Answer by stanbon(75887)

3/7 x (5/2) =
The parentheses helps clarify the meaning of the problem
(3/7)(5/2) = 15/14
=========================
Cheers,
Stan H.